# Fedor V. Fomin

### v1

@@ 4 Conclusion
We conclude with several open research questions. All our results concern undirected graphs. The complexity of all these problems on directed graphs of small independence number remains open. For example, we do not know whether Hamiltonian Cycle is $\operatorname{{\sf NP}}$-complete on directed graphs with an independence number $k$ for some fixed $k\geq 2$. But we cannot exclude that the problem may be in $\operatorname{{\sf X}P}$ or even $\operatorname{{\sf FPT}}$ parameterized by $k$. Parameterized complexity offers a powerful toolbox for obtaining lower bounds for structural parameterization [${}^{+}$11, ${}^{+}$19, FGLS10, LMS18]. However, all these reductions use gadgets with large independence numbers. Obtaining algorithmic lower bounds for parameterization with small independence number appears more challenging.

### v2

@@ 5 Conclusion
We conclude with several open research questions. All our results concern undirected graphs. The complexity of all these problems on directed graphs of small independence number remains open. For example, we do not know whether Hamiltonian Cycle is $\operatorname{{\sf NP}}$-complete on directed graphs with an independence number $k$ for some fixed $k\geq 2$. But we cannot exclude that the problem may be in $\operatorname{{\sf X}P}$ or even $\operatorname{{\sf FPT}}$ parameterized by $k$. Parameterized complexity offers a powerful toolbox for obtaining lower bounds for structural parameterization [CNP+11, FGL+19, FGLS10, LMS18]. However, all these reductions use gadgets with large independence numbers. Obtaining algorithmic lower bounds for parameterization with small independence number appears more challenging.